OpenStudy (anonymous):

3. Having survived the meteor impact, thanks to some last-minute evasive maneuvers, the Mathonauts now set their sights on their Interstellar Headquarters. The Interstellar Headquarters orbits the Earth based on the equation y2 + x2 = 40,000. Using the original trajectory of the ship and complete sentences, explain to the pilot how to find where the ship’s path will cross the Interstellar Headquarters’ path. The equation I chose was y=2x+5 and I need to know how to plug it into y and solve for x.

jimthompson5910 (jim_thompson5910):

you have \[\Large y^2 + x^2 = 40,000\]

jimthompson5910 (jim_thompson5910):

replace y with 2x+5. This is valid because y = 2x+5 \[\Large y^2 + x^2 = 40,000\] \[\Large (2x+5)^2 + x^2 = 40,000\]

OpenStudy (anonymous):

I know I need to get x to one side and solve

jimthompson5910 (jim_thompson5910):

then you expand out (2x+5)^2 \[\Large (2x+5)^2 + x^2 = 40,000\] \[\Large (2x+5)(2x+5) + x^2 = 40,000\] \[\Large 2x(2x+5)+5(2x+5) + x^2 = 40,000\] \[\Large 4x^2+10x+10x+25 + x^2 = 40,000\]

jimthompson5910 (jim_thompson5910):

I'll let you finish up

OpenStudy (anonymous):

Do I have to foil the x^2 so I can get x by itself and to one side of the equation?

jimthompson5910 (jim_thompson5910):

I recommend combining like terms for the next step

OpenStudy (anonymous):

I was thinking of that but I have to get x to one side of the equation because I am trying to solve for x. I have to find two points so I can put them on a graph when I am done.

jimthompson5910 (jim_thompson5910):

you'll need to use the quadratic formula

jimthompson5910 (jim_thompson5910):

but first you need to get everything to one side

OpenStudy (anonymous):

could you help me with that? I just need help with plugging in my equation for y and getting x to one side of the equation and then I can solve the rest on my own.

jimthompson5910 (jim_thompson5910):

\[\Large 4x^2+10x+10x+25 + x^2 = 40,000\] \[\Large 4x^2+10x+10x+25 + x^2-40,000 = 0\] \[\Large 5x^2+20x-39,975 = 0\]

jimthompson5910 (jim_thompson5910):

what's the next step?

OpenStudy (anonymous):

What happened to the y?

jimthompson5910 (jim_thompson5910):

I replaced it with 2x+5

jimthompson5910 (jim_thompson5910):

since y = 2x+5

jimthompson5910 (jim_thompson5910):

I did that up near the top

OpenStudy (anonymous):

Im not sure from here because x isnt on one side

jimthompson5910 (jim_thompson5910):

do you know the quadratic formula?

OpenStudy (anonymous):

ax2 + bx + c = 0

jimthompson5910 (jim_thompson5910):

that's part of it

jimthompson5910 (jim_thompson5910):

what was that you posted?

OpenStudy (anonymous):

i was trying to post the second part but it came out funny. it was x=-b plus or negative square root b^2 -4ac/2a

OpenStudy (anonymous):

I really need help Im not sure how to solve it until only x is on one side and not 0.

jimthompson5910 (jim_thompson5910):

I gotcha, so this \[\Large x = \frac{-b\pm \sqrt{b^2-4ac}}{2a}\]

jimthompson5910 (jim_thompson5910):

what are a,b,c in this case?

OpenStudy (anonymous):

a=5 b=20 and c= -39,975

jimthompson5910 (jim_thompson5910):

so you'll plug those values into the formula I wrote above

OpenStudy (anonymous):

Do I keep the negative on 39,975

jimthompson5910 (jim_thompson5910):

yes you do, c is negative

OpenStudy (anonymous):

This is what I came up with. x=87.4371287553441 x=−91.4371287553441

jimthompson5910 (jim_thompson5910):

let me check

jimthompson5910 (jim_thompson5910):

I'm getting the same. Those are the approximate solutions

OpenStudy (anonymous):

Is there something easier we could do? because this is too much work for a simple word problem. I only need one x solution and one y solution. It should be where I plug the equation for y move x to one side of the equation by subtracting or something and then solve and whatever my answer is I plug it back in for y and solve. I need something easy.

jimthompson5910 (jim_thompson5910):

with quadratics, you could either factor or use the quadratic formula or complete the square the quadratic formula works with any quadratic while factoring only works with some quadratics

OpenStudy (anonymous):

I know that but this is not like my homework lesson. in the lesson we solved by substitution or elimination.

OpenStudy (anonymous):

It was way easier than using the quadratic formula

jimthompson5910 (jim_thompson5910):

well in that first step I used substitution to replace y with 2x+5

OpenStudy (anonymous):

after plugging in the equation for y could we just factor out x and then subtract x to one side?

jimthompson5910 (jim_thompson5910):

that method won't work unfortunately

OpenStudy (anonymous):

oh... well then what am I supposed to do with two x solutions?

OpenStudy (anonymous):

Im only supposed to have one x and one y solutions to plot on a graph

OpenStudy (anonymous):

the answer at the end of all this is supposed to be where these two equations intersect

jimthompson5910 (jim_thompson5910):

does it say that x has to be positive?

OpenStudy (anonymous):

no but at this point i think it is supposed to be positive

OpenStudy (anonymous):

You can substitute your original trajectory ( y=5x+1 ) into the equation that the Interstellar Headquarters is based on ( y2+x2=40,000 ). You will be with the equation 2( 5x+1 )+x2 = 40,000 and you will solve for X. After solving for X, you will have X=3333.12 for your value. You will now take this value back into one of your original equations. You will now have the equation y=5 ( 3333.17 )+1 and will solve for Y. Your value for Y will be Y=16666.85. You now know the ship’s path will cross the Interstellar Headquarters’ at the coordinate ( 3333.17, 16666.85 ).

OpenStudy (anonymous):

This is from someone with the same question as me and they made it real simple just ignore all the question marks

jimthompson5910 (jim_thompson5910):

they made an error

jimthompson5910 (jim_thompson5910):

they thought the y^2 was 2y

OpenStudy (anonymous):

maybe it is supposed to be that because if its y^2 and x^2 then we would come up with more than one x solution and i can only have one.

jimthompson5910 (jim_thompson5910):

well usually y2 is often y^2 because I've seen where the symbols don't paste properly

OpenStudy (anonymous):

could we try solving it the same way that person did and see what we come up with?

jimthompson5910 (jim_thompson5910):

I solved their equation and I got 2 solutions also

jimthompson5910 (jim_thompson5910):

not just 1 solution

jimthompson5910 (jim_thompson5910):

so I'm not sure which solution they want

OpenStudy (anonymous):

maybe if we solved for y we might come up with 2 solutions as well

jimthompson5910 (jim_thompson5910):

yeah I think that will happen as well